# Flow Matching at Scale: A Machine Learning Framework for Efficient Large-Size Sampling of Many-Body Systems

### v3

@@ V Result II: Reliable Extrapolation of System Sizes
A key strength of our Flow Matching approach is its ability to extrapolate to different system sizes without retraining. This capability is central to our study’s primary goal: to demonstrate that Flow Matching can efficiently capture the essential underlying physics of a system from limited data, prioritizing qualitative consistency over precise quantitative reproduction. As shown in Fig. 5, a single model trained exclusively on $32\times 32$ configurations successfully generates samples for a $64\times 64$ system. The results for key observables exhibit strong qualitative agreement with the reference MC data, confirming that the learned physical rules are generalized effectively.
@@
Figure 7: Susceptibility $\chi(T)$ for various system sizes ($L=16,64,128$), demonstrating the size-extrapolation capability of a single Flow Matching (FM) model. The model was trained exclusively on $32\times 32$ data ($T\in[0.1,2.0]$, $\Delta T=0.01$, 1500 configs/T). Although quantitative deviations from the MCMC (MC) reference increase as the system size diverges from the training size, the FM model qualitatively captures the essential physics. Specifically, it correctly reproduces the characteristic shift and sharpening of the susceptibility peak with system size, a key signature of the BKT transition.
This extrapolation is robust across a wide range of system dimensions. As shown in Figs. 6 and 7, the same model trained on $32\times 32$ generates configurations for lattices as small as $16\times 16$ and as large as $128\times 128$. Although quantitative deviations increase with the extrapolation range, the model correctly captures essential finite-size effects, such as the characteristic shift and sharpening of the susceptibility peak, a key signature of the BKT transition. This preservation of qualitative features underscores the model’s ability to learn fundamental physical behavior.
@@
Figure 8: Finite-size scaling of the BKT transition. The susceptibility peak temperature $T_{\text{peak}}(L)$ is plotted against $(\ln L)^{-2}$ for various lattice sizes (L=16, 20, 24, 28, 32, 40, 48, 64, 80, 96, 128L=16,\;20,\;24,\;28,\;32,\;40,\;48,\;64,\;80,\;96,\;128). All data points were generated from a single Flow Matching model trained exclusively on $32\times 32$ data ($T\in[0.1,2.0]$, $\Delta T=0.01$, 1500 configs/T). The linear trend, when extrapolated to the thermodynamic limit ($L\to\infty$), yields $T_{\text{BKT}}\approx 0.895$, in good agreement with the established value of $\approx 0.8935$ [31]. This result confirms that the model has successfully captured the underlying scale-invariant physics.
@@
A finite-size scaling analysis tests the model’s ability to learn scale-invariant physics. Our primary objective here is to qualitatively demonstrate this emergent capability, rather than to achieve a high-precision estimate of $T_{BKT}$. We acknowledge that spin stiffness is a more rigorous metric for analyzing the BKT transition [31]. However, as discussed in our error analysis (Sec. VI), the model’s predictions for fluctuation-related quantities are inherently less precise, a characteristic compounded by our deliberate use of a limited, low-precision training dataset (1500 configurations per temperature from MCMC). Consequently, the statistical noise in the generated stiffness data, particularly near the critical point, was too significant to permit a reliable scaling analysis. We transparently document this behavior alongside other observables in Appendix C.
@@
In contrast, the location of the susceptibility peak, $T_{\text{peak}}(L)$, provided a more robust and clearly identifiable feature for demonstrating the expected scaling behavior. We therefore adopted this metric for our analysis. Using the single model trained on $32\times 32$ data, we located $T_{\text{peak}}(L)$ for various lattice sizes ($L=16$ to $128$). As depicted in Fig. 8, a plot of $T_{\text{peak}}(L)$ versus $(\ln L)^{-2}$ reveals the expected linear relationship [32]. Extrapolating to the thermodynamic limit ($L\to\infty$) yields an estimated $T_{\text{BKT}}\approx 0.895$, in reasonable agreement with the established value of $\approx 0.8935$ [31]. This result, achieved without high-precision training data or retraining, confirms that the model has successfully learned the size-independent physical principles of the XY model, highlighting its efficiency for studying critical phenomena.
VI Result III: Errors Analysis
@@
To quantitatively assess the performance of our Flow Matching model, we compute the mean squared error (MSE) between its predictions and the ground-truth MCMC data for key physical observables, as defined in Eq. (21), where TT is temperature and $N_{T}$ is the number of temperature points. Fig. 9 presents this error analysis for a model trained on $32\times 32$ lattices and then used to generate samples for systems of size L=16, 32, 64,L=16,\;32,\;64, and $128$.
@@
Figure 9: Mean squared error (MSE) as a function of temperature, quantifying the error of the size extrapolation method. The MSE is calculated between the Flow Matching and reference MCMC results for several observables: susceptibility $\chi$, spin stiffness $\rho_{s}$, energy $E$, magnetization $m$, and vortex density $\rho_{v}$. A single FM model, trained on a $32\times 32$ lattice ($T\in[0.1,2.0]$, $\Delta T=0.01$), was used to generate data for systems of size L=16, 32, 64,L=16,\;32,\;64, and $128$. The evaluation for L=16, 32L=16,\;32 and $64$ uses the same temperature grid as the training, while the evaluation for the largest system, $L=128$, is performed on a finer grid focused on the critical region ($T\in[0.8,1.05]$, $\Delta T=0.005$).
The analysis reveals two important trends. First, the errors for most of quantities are remarkably small, underscoring the high fidelity of the generated samples even when extrapolating to system sizes four times larger than the training data. The errors are small for the $L=16$ system, demonstrating robust interpolation to smaller sizes, and predictably increase with the extrapolation range to $L=64$ and $L=128$. Second, and more revealingly, the model exhibits higher accuracy for first-moment quantities like energy ($E$) and magnetization ($M$) compared to second-moment quantities related to fluctuations, such as susceptibility ($\chi$) and spin stiffness ($\rho_{s}$). This discrepancy is most pronounced near the critical temperature (see Fig. 7), where physical fluctuations are maximal.
@@
To minimize this total error, the model must balance its variance (due to finite training data) against a systematic bias towards the mean of the true conditional velocity field. This optimization naturally produces a learned vector field that under-represents the actual variance of the physical process. Furthermore, this initial inaccuracy in the learned field’s variance is compounded during the inference stage. Sampling requires numerical integration of the probability flow ODE with a finite number of steps. As suggested in Fig. 2, the vector guiding the trajectory in each discrete step of the numerical solver is an approximation. Minor errors in the velocity field, particularly regarding its variance, accumulate during the iterative integration process. This can cause the final generated configurations to incorrectly capture the actual spread of the target distribution, leading to a larger error in the variance of the resulting ensemble. Consequently, while the generated samples accurately reflect first-moment quantities such as energy and magnetization, the fluctuation-related properties of the generated ensemble are less reliable. Despite FM-generated samples showing larger errors in fluctuation-related physical statistics, which grow more significantly with size extrapolation, the model remains qualitatively accurate enough to capture essential physics, such as the characteristic peaks in susceptibility, making it a useful tool for identifying phase transitions and performing finite-size scaling analysis.
@@
With natural logarithms, $0\leq\operatorname{JSD}(\mathcal{P}\|\mathcal{Q})\leq\ln 2$, the upper bound $\ln 2$ is reached precisely when $\mathcal{P}$ and $\mathcal{Q}$ are completely unrelated. Fig. 10 shows the JSD of energy distribution at different temperatures. Across most of the temperature range, including the phase transition point ($T\approx 1.0$ for system size $32\times 32$), the JSD is less than 0.1, indicating that the Flow Matching sampling results agree with the Monte Carlo simulation. Only near zero temperature does the JSD exceed 0.1. The histograms embedded in Fig. 10 show that the energy distribution becomes increasingly concentrated as the temperature decreases (the bar widths in all histograms are the same). Because Flow Matching uses a neural network model that is simultaneously trained and sampled over the entire temperature range, the figure reveals that under the same conditions, Flow Matching performs relatively poorly for extremely narrow, concentrated distributions.
@@
Figure 10: Jensen–Shannon Divergence (JSD) between the energy distributions generated by the Flow Matching (FM) and MCMC methods as a function of temperature. The analysis was performed on a $32\times 32$ lattice, using an FM model trained and evaluated under the specified conditions ($T\in[0.1,2.0]$, $\Delta T=0.01$, 1500 configs/T). The insets show the corresponding energy histograms at four representative temperatures. All histograms are plotted with identical bin widths to compare the distributions’ spread visually. The JSD is highest at very low temperatures, which, as illustrated in the T=0.1 inset, corresponds to a highly concentrated energy distribution that is more challenging for the model to reproduce precisely.
VII Result IV: Computational Efficiency for Large-Scale Studies
@@
The primary computational advantage of our Flow Matching approach stems from its ability to generalize across system sizes, a capability traditional MCMC methods lack. We compare the total time cost required for a typical finite-size scaling analysis to quantify the practical implications of this paradigm shift in workflow. For our FM model, this involves a one-time training cost on a small lattice ($32\times 32$), followed by rapid inference on a series of larger lattices. The MCMC method requires running a separate, full simulation for each lattice size. We use our highly optimized, GPU-accelerated MCMC algorithm (see Appendix F) as the baseline for this comparison. All timing benchmarks were performed on a single workstation; detailed hardware and software specifications are provided in Appendix G. The key metric for this evaluation is the Sampling Cost, defined as the average wall-clock time to generate a single statistically independent sample:
@@
Figure 11: Computational cost per statistically independent sample as a function of system size $L$. The figure compares the performance of our GPU-accelerated MCMC method against the Flow Matching (FM) model, which was trained only on the $L=32$ system. The FM cost is presented in two ways: the inference cost alone and the total cost including the amortized one-time training expense, as defined in Eq. (26). The results show a crossover in efficiency, with the FM approach becoming significantly faster than MCMC for large system sizes, establishing its advantage for large-scale studies.
As shown in Fig. 11, while the initial training constitutes a significant upfront investment, this cost is amortized over all subsequent sample generation for various larger system sizes. The crossover point at $L=128$ demonstrates that our method offers a substantial net reduction in computational resources for comprehensive studies of critical phenomena approaching the thermodynamic limit. This highlights that the actual efficiency of our model lies not in single-sample speed, but in its ability to circumvent the need for repeated, costly simulations across different scales.
VIII Physical Interpretation for The Size Extrapolation
The model’s remarkable ability to generalize across system sizes stems from a carefully engineered synergy between the learning framework, the model architecture, and the fundamental nature of the physical system. Unlike methods that learn size-specific global statistics, our approach is designed to distill the underlying, scale-invariant local rules dictated by the Hamiltonian. The following subsections will elaborate on the key components that enable this capability. We will first discuss how the strong inductive biases of Convolutional Neural Networks (CNNs), namely locality and translation equivariance, are naturally aligned with the structure of local physical interactions. We will then explain how the specific U-Net architecture leverages these biases across multiple scales to capture both local details and emergent long-range correlations. Finally, we will synthesize these points to argue that the model learns a representation of the local physics itself, which is inherently size-agnostic and thus powerfully generalizable.
@@ VIII.1 Inductive Bias
Inductive bias refers to implicit assumptions within machine learning algorithms that enable generalization to unseen data. The effectiveness of an algorithm depends significantly on how well its inductive biases align with the characteristics of the given task. Thus, selecting an appropriate model involves matching these biases to the specific nature of the problem. For example, several recent studies similar to this work have pointed out that the equivariance of neural networks (a kind of inductive bias) has significant advantages in dealing with physical systems with symmetries [33, 34].
@@ VIII.2 U-Net Structure
While the inductive biases of locality and translation equivariance make standard CNNs well suited to systems with local interactions, they struggle with long-range correlations near criticality because the receptive field grows only linearly with depth of network. The U-Net architecture [16], built by the encoder and decoder structure, mitigates this by efficiently integrating information across multiple spatial scales, especially when paired with information-preserving resizing (e.g., pixel-shuffle–based rearrangements (see Fig. 12[35]). Crucially, even when operating on downsampled feature maps in the encoder that summarize large-scale structure, convolution remains a local operator: kernels still act on neighboring sites of the current lattice, thereby learning transformation rules that are consistent across resolutions—i.e., local physical laws expressed in renormalized coordinates that remain (approximately) scale-invariant. To ensure that the model accurately reflects the underlying physics, we align its architectural choices with the simulation’s boundary conditions. All Monte Carlo simulations in this work employ periodic boundary conditions. Correspondingly, all convolutional layers within the U-Net architecture are implemented with circular padding. This crucial detail enforces the same periodic topology on the feature maps as on the physical lattice, ensuring that the learned local rules are consistent with the system’s global structure.
Figure 12: The pixel shuffle operation [35], a lossless resizing technique used in our U-Net architecture. The encoder employs a space-to-depth pixel shuffle (illustrated from left to right) to downsample spatial resolution while increasing channel depth. Conversely, the decoder uses an inverse depth-to-space pixel shuffle (right to left) for upsampling to reconstruct the spatial dimensions.
Mechanistically, the encoder uses an pixel shuffle (space-to-depth) [35] to reduce spatial resolution without discarding information, mapping, for example, each $2\times 2$ block into four channels (see Fig. 12). This preserves translation equivariance and expands the effective receptive field exponentially with depth [36], while convolutions at coarser resolutions still implement local updates on block variables. The decoder applies the inverse pixel shuffle (depth-to-space) to reconstruct high-resolution outputs (see Fig. 12), and skip connections concatenate encoder features with aligned decoder features. This fusion preserves fine-grained locality while enforcing cross-scale consistency, allowing the network to integrate global context without sacrificing the learning of scale-consistent local rules at each resolution. For more architectural and technical details, see the Appendix D.
@@
To empirically validate that the U-Net’s inductive biases are key to its successful size extrapolation, we conducted a comparative study against a Diffusion Transformer (DiT) model [37], which lacks strong, built-in locality. The U-Net significantly outperformed the DiT, especially in the extrapolation task, despite having substantially fewer parameters. This result supports our claim that architectural alignment with the problem’s local nature is critical. The detailed experimental setup, results, and discussion of this comparison are presented in Appendix E.
VIII.3 Learning Local Physics Instead of Global Statistics
The MCMC simulation is inherently tied to the global parameters of the system it simulates. It learns nothing beyond the equilibrium properties of the size and temperature of that single system. The process is a powerful tool for the analysis of a fixed system, not for learning generalizable physical principles across scales. In contrast, Flow Matching operates on a different principle. It does not sample a stationary distribution but instead learns a deterministic, non-equilibrium transformation. The goal is to construct a continuous path that transforms a simple, unstructured probability distribution (e.g., Gaussian noise) into the complex target data distribution (in our case, the physical spin configurations from MCMC). The characteristics of the Flow Matching Markov process are:

### v4

@@ V Result II: Extrapolation for Finite-Size Scaling Analysis
A key strength of our Flow Matching approach is its ability to extrapolate to different system sizes without retraining. This capability is central to our study’s primary goal: to demonstrate that Flow Matching can efficiently capture the critical scaling behavior of a system from limited data, prioritizing the identification of phase boundaries over precise quantitative reproduction. As shown in Fig. 5, a single model trained exclusively on $32\times 32$ configurations successfully generates samples for a $64\times 64$ system. The results for key observables exhibit consistent trends with the reference MC data, confirming that the learned local interaction patterns are generalized effectively.
@@
Figure 7: Susceptibility $\chi(T)$ for various system sizes ($L=16,64,128$), demonstrating the size-extrapolation capability of a single Flow Matching (FM) model. The model was trained exclusively on $32\times 32$ data ($T\in[0.1,2.0]$, $\Delta T=0.01$, 1500 configs/T). Although quantitative deviations from the MCMC (MC) reference increase as the system size diverges from the training size, the FM model qualitatively captures the essential physics. Specifically, it correctly reproduces the characteristic shift and sharpening of the susceptibility peak with system size, a key signature of the BKT transition. This property allows us to train with only small-sized Monte Carlo data and directly generate high-quality, larger-sized initial spin configurations, which is beneficial for subsequent large-sized Monte Carlo simulations and greatly improves simulation efficiency.
This extrapolation is robust across a wide range of system dimensions. As shown in Figs. 6 and 7, the same model trained on $32\times 32$ generates configurations for lattices as small as $16\times 16$ and as large as $128\times 128$. It is crucial to note that the model was never exposed to the global statistics of the larger systems during training. The emergence of size-dependent features, such as the shifting susceptibility peak, arises because the model applies learned local interaction patterns consistently across the larger lattice geometry. Although quantitative deviations increase with the extrapolation range (likely due to the accumulation of integration errors and the lack of exact long-range constraints in the training data), the model successfully reproduces the essential finite-size effects and the characteristic BKT scaling signatures.
@@
Figure 8: Finite-size scaling of the BKT transition. The susceptibility peak temperature $T_{\text{peak}}(L)$ is plotted against $(\ln L)^{-2}$ for various lattice sizes (L=16, 20, 24, 28, 32, 40, 48, 64, 80, 96, 128L=16,\;20,\;24,\;28,\;32,\;40,\;48,\;64,\;80,\;96,\;128). All data points were generated from a single Flow Matching model trained exclusively on $32\times 32$ data ($T\in[0.1,2.0]$, $\Delta T=0.01$, 1500 configs/T). The linear trend, when extrapolated to the thermodynamic limit ($L\to\infty$), yields $T_{\text{BKT}}\approx 0.907$, which is very close to the established value of $\approx 0.8935$ [Hsieh_2013_Finite_size_scaling]. This result confirms that the model has successfully captured the underlying scale-invariant physics. Note that it is neither possible nor necessary to plot the Monte Carlo simulation results here for comparison, since we are using Monte Carlo data of a single size for training.
@@
A finite-size scaling analysis tests the model’s ability to learn scale-invariant physics. Our primary objective here is to qualitatively demonstrate this emergent capability, rather than to achieve a high-precision estimate of $T_{BKT}$. We acknowledge that spin stiffness is a more rigorous metric for analyzing the BKT transition [Hsieh_2013_Finite_size_scaling]. However, as discussed in our error analysis (Sec. VI), the model’s predictions for fluctuation-related quantities are inherently less precise, a characteristic compounded by our deliberate use of a limited, low-precision training dataset (1500 configurations per temperature from MCMC). Consequently, the statistical noise in the generated stiffness data, particularly near the critical point, was too significant to permit a reliable scaling analysis. We transparently document this behavior alongside other observables in Appendix C.
@@
In contrast, the location of the susceptibility peak, $T_{\text{peak}}(L)$, provided a more robust and clearly identifiable feature for demonstrating the expected scaling behavior. We therefore adopted this metric for our analysis. Using the single model trained on $32\times 32$ data, we located $T_{\text{peak}}(L)$ for various lattice sizes ($L=16$ to $128$). As depicted in Fig. 8, a plot of $T_{\text{peak}}(L)$ versus $(\ln L)^{-2}$ reveals the expected linear relationship [PhysRevB.65.184405]. Extrapolating to the thermodynamic limit ($L\to\infty$) yields an estimated $T_{\text{BKT}}\approx 0.907$, in reasonable agreement with the established value of $\approx 0.8935$ [Hsieh_2013_Finite_size_scaling]. This result, achieved without high-precision training data or retraining, suggests that the model has successfully learned the scale-invariant correlation patterns necessary for locating the critical point of the XY model, highlighting its efficiency for studying critical phenomena.
VI Result III: Deviation Analysis and Limits of Estimation
@@
To quantitatively assess the performance of our Flow Matching model, we compute the mean squared error (MSE) between its generated estimates and the reference MCMC data for key physical observables, as defined in Eq. (22), where TT is temperature and $N_{T}$ is the number of temperature points. Fig. 9 presents this error analysis for a model trained on $32\times 32$ lattices and then used to generate samples for systems of size L=16, 32, 64,L=16,\;32,\;64, and $128$.
@@
Figure 9: Mean squared error (MSE) as a function of temperature, quantifying the deviation of the size extrapolation method. The MSE is calculated between the Flow Matching and reference MCMC results for several observables: susceptibility $\chi$, spin stiffness $\rho_{s}$, energy $E$, magnetization $m$, and vortex density $\rho_{v}$. A single FM model, trained on a $32\times 32$ lattice ($T\in[0.1,2.0]$, $\Delta T=0.01$), was used to generate data for systems of size L=16, 32, 64,L=16,\;32,\;64, and $128$. The evaluation for L=16, 32L=16,\;32 and $64$ uses the same temperature grid as the training, while the evaluation for the largest system, $L=128$, is performed on a finer grid focused on the critical region ($T\in[0.8,1.05]$, $\Delta T=0.005$).
The analysis reveals two important trends. First, the deviations for first-moment quantities remain minimal, indicating that the generated configurations retain sufficient structural information even when extrapolating to system sizes four times larger than the training data. The errors are small for the $L=16$ system, demonstrating robust interpolation to smaller sizes, and predictably increase with the extrapolation range to $L=64$ and $L=128$. Second, and more revealingly, the model exhibits higher consistency for first-moment quantities like energy ($E$) and magnetization ($M$) compared to second-moment quantities related to fluctuations, such as susceptibility ($\chi$) and spin stiffness ($\rho_{s}$). This discrepancy is most pronounced near the critical temperature (see Fig. 7), where critical fluctuations dominate.
@@
To minimize this total error, the model must balance its variance (due to finite training data) against a systematic bias towards the mean of the true conditional velocity field. This optimization naturally produces a learned vector field that prioritizes the mean field trend over the precise variance of the target distribution. Furthermore, this initial inaccuracy in the learned field’s variance is compounded during the inference stage. Sampling requires numerical integration of the probability flow ODE with a finite number of steps. As suggested in Fig. 2, the vector guiding the trajectory in each discrete step of the numerical solver is an approximation. Minor errors in the velocity field, particularly regarding its variance, accumulate during the iterative integration process. This can cause the final generated configurations to underestimate the actual spread of the target distribution, leading to a larger error in the variance of the resulting ensemble. Consequently, while the generated samples accurately reflect first-moment quantities such as energy and magnetization, the fluctuation-related properties of the generated ensemble are less reliable. Despite FM-generated samples showing larger errors in fluctuation-related physical statistics, which grow more significantly with size extrapolation, the model remains robust in identifying the temperature location of characteristic peaks in susceptibility. This makes it a highly efficient tool for locating phase transitions and performing finite-size scaling analysis, even if the absolute magnitude of fluctuations is underestimated.
@@
With natural logarithms, $0\leq\operatorname{JSD}(\mathcal{P}\|\mathcal{Q})\leq\ln 2$, the upper bound $\ln 2$ is reached precisely when $\mathcal{P}$ and $\mathcal{Q}$ are completely unrelated. Fig. 10 shows the JSD of energy distribution at different temperatures. Across most of the temperature range, including the phase transition point ($T\approx 1.0$ for system size $32\times 32$), the JSD is less than 0.1, indicating that the generated energy distributions are statistically consistent with the Monte Carlo reference. Only near zero temperature does the JSD exceed 0.1. The histograms embedded in Fig. 10 show that the energy distribution becomes increasingly concentrated as the temperature decreases (the bar widths in all histograms are the same). Because Flow Matching uses only one neural network model that is simultaneously trained and sampled over the entire temperature range, the figure reveals that under the same conditions, the generative model faces challenges in resolving the fine-grained variance of extremely narrow distributions. The bias of distributions near the phase transition point also indicates that the neural network is still some distance from generating distributions that are entirely consistent with physics. However, it is still sufficient to quickly locate the phase transition point and generate a high-quality initial spin configuration for Monte Carlo simulation.
@@
Figure 10: Jensen–Shannon Divergence (JSD) between the energy distributions generated by the Flow Matching (FM) and MCMC methods as a function of temperature. The analysis was performed on a $32\times 32$ lattice, using an FM model trained and evaluated under the specified conditions ($T\in[0.1,2.0]$, $\Delta T=0.01$, 1500 configs/T). The insets show the corresponding energy histograms at four representative temperatures. All histograms are plotted with identical bin widths to compare the distributions’ spread visually. As can be seen in the figure, JSD has a significant value near zero temperature ($T=0.1$) and the phase transition point ($T\approx 0.89$), which means that the neural network still cannot perform perfectly in the extreme distribution region, but it is sufficient for quickly locating the phase transition point and generating high-quality initial spin configurations for Monte Carlo simulation.
VII Result IV: Efficiency in Multi-Scale Critical Analysis
@@
The primary computational advantage of our Flow Matching approach stems from its ability to generalize across system sizes, allowing for rapid estimation without re-simulation. We compare the total time cost required for a typical finite-size scaling analysis to quantify the practical implications of this methodological approach. For our FM model, this involves a one-time training cost on a small lattice ($32\times 32$), followed by rapid inference on a series of larger lattices. The MCMC method requires running a separate, full simulation for each lattice size. We use our highly optimized, GPU-accelerated MCMC algorithm (see Appendix F) as the baseline for this comparison. All timing benchmarks were performed on a single workstation; detailed hardware and software specifications are provided in Appendix G. The key metric for this evaluation is the Sampling Cost, defined as the average wall-clock time to generate a single independent configuration:
@@
Figure 11: Computational cost per independent configuration as a function of system size $L$. The figure compares the performance of our GPU-accelerated MCMC method against the Flow Matching (FM) model, which was trained only on the $L=32$ system. The FM cost is presented in two ways: the inference cost alone and the total cost including the amortized one-time training expense, as defined in Eq. (27). The results show a crossover in efficiency, which can be used to locate phase transition points across multiple sizes quickly or to rapidly generate high-quality initial spin configurations for large-scale Monte Carlo simulations. Because the cost of training is a one-time expense. This method reduces the burden of traditional Monte Carlo simulations by leveraging the ability of trained neural networks to be reused or fine-tuned for other tasks.
As shown in Fig. 11, while the initial training constitutes a significant upfront investment, this cost is amortized over all subsequent sample generation for various larger system sizes. The crossover point at $L=128$ demonstrates that our method offers a streamlined workflow for locating critical points in the thermodynamic limit. This highlights that the actual efficiency of our model lies not in single-sample speed, but in its ability to circumvent the need for repeated, costly simulations across different scales.
VIII Physical Interpretation of Size Extrapolation
The model’s remarkable ability to generalize across system sizes stems from a carefully engineered synergy between the learning framework, the model architecture, and the fundamental nature of the physical system. Unlike methods that learn size-specific global statistics, our approach is designed to distill the underlying, scale-invariant local correlation patterns dictated by the Hamiltonian. The following subsections will elaborate on the key components that enable this capability. We will first discuss how the strong inductive biases of Convolutional Neural Networks (CNNs), namely locality and translation equivariance, are naturally aligned with the structure of local physical interactions. We will then explain how the specific U-Net architecture leverages these biases across multiple scales to capture both local details and emergent long-range correlations. Finally, we will synthesize these points to argue that the model learns a representation of the local generative dynamics for locating critical points.
@@ VIII.1 Inductive Bias
Inductive bias refers to implicit assumptions within machine learning algorithms that enable generalization to unseen data. The effectiveness of an algorithm depends significantly on how well its inductive biases align with the characteristics of the given task. Thus, selecting an appropriate model involves matching these biases to the specific nature of the problem. For example, several recent studies similar to this work have pointed out that the equivariance of neural networks (a kind of inductive bias) has significant advantages in dealing with physical systems with symmetries [Schuh:2025fky_Hubbard_Model_NF, Equivariant_flow_matching].
@@ VIII.2 U-Net Structure
While the inductive biases of locality and translation equivariance make standard CNNs well suited to systems with local interactions, they struggle with long-range correlations near criticality because the receptive field grows only linearly with depth of network. The U-Net architecture [ronneberger2015unet], built by the encoder and decoder structure, mitigates this by efficiently integrating information across multiple spatial scales, especially when paired with information-preserving resizing (e.g., pixel-shuffle–based rearrangements (see Fig. 12[shi2016pixel_shuffle]). Crucially, even when operating on downsampled feature maps in the encoder that summarize large-scale structure, convolution remains a local operator: kernels still act on neighboring sites of the current lattice, thereby learning transformation rules that are consistent across resolutions—i.e., effective local interactions in renormalized coordinates that remain (approximately) scale-invariant. To ensure that the model accurately reflects the underlying physics, we align its architectural choices with the simulation’s boundary conditions. All Monte Carlo simulations in this work employ periodic boundary conditions. Correspondingly, all convolutional layers within the U-Net architecture are implemented with circular padding. This crucial detail enforces the same periodic topology on the feature maps as on the physical lattice, ensuring that the learned local rules are consistent with the system’s global structure.
Figure 12: The pixel shuffle operation [shi2016pixel_shuffle], a lossless resizing technique used in our U-Net architecture. The encoder employs a space-to-depth pixel shuffle (illustrated from left to right) to downsample spatial resolution while increasing channel depth. Conversely, the decoder uses an inverse depth-to-space pixel shuffle (right to left) for upsampling to reconstruct the spatial dimensions.
Mechanistically, the encoder uses an pixel shuffle (space-to-depth) [shi2016pixel_shuffle] to reduce spatial resolution without discarding information, mapping, for example, each $2\times 2$ block into four channels (see Fig. 12). This preserves translation equivariance and expands the effective receptive field exponentially with depth [wang2024theoreticalanalysis_inductive_biases], while convolutions at coarser resolutions still implement local updates on block variables. The decoder applies the inverse pixel shuffle (depth-to-space) to reconstruct high-resolution outputs (see Fig. 12), and skip connections concatenate encoder features with aligned decoder features. This fusion preserves fine-grained locality while enforcing cross-scale consistency, allowing the network to integrate global context without sacrificing the learning of scale-consistent local rules at each resolution. For more architectural and technical details, see the Appendix D.
@@
To empirically validate that the U-Net’s inductive biases are key to its successful size extrapolation, we conducted a comparative study against a Diffusion Transformer (DiT) model [peebles2023DiT], which lacks strong, built-in locality. The U-Net significantly outperformed the DiT, especially in the extrapolation task, despite having substantially fewer parameters. This result supports our claim that architectural alignment with the problem’s local nature is critical. The detailed experimental setup, results, and discussion of this comparison are presented in Appendix E.
VIII.3 Capturing Local Dynamics for Scalable Generation
The MCMC simulation is inherently tied to the global parameters of the system it simulates. It learns nothing beyond the equilibrium properties of the size and temperature of that single system. The process is a powerful tool for the analysis of a fixed system, not for learning generalizable interaction patterns across scales. In contrast, Flow Matching operates on a different principle. It does not sample a stationary distribution but instead learns a deterministic, non-equilibrium transformation. The goal is to construct a continuous path that transforms a simple, unstructured probability distribution (e.g., Gaussian noise) into the complex target data distribution (in our case, the physical spin configurations from MCMC). The characteristics of this generative framework include: